Solve the equation by using the LCD.
4/x +1=2x+2/x
(4 over x + 1 =2x+2 over x)

2x/x+2 -4=6/x
(2x over x+2 -4 =6 over x)
For some reason I can't figure these out. I've tried and they just aren't making sense to me

Answer by stanbon(75887)

Solve the equation by using the LCD.
4/(x+1) = (2x+2)x
4/(x+1) = 2(x+1)/x
--------------------------
LCD:: x(x+1)
Multiply both sides by the LCD to get:
4x = 2(x+1)^2
2x = (x+1)^2
x^2 + 2x +1 = 2x
x^2 + 1 = 0
x is either +-i or there is no Real Number solution.
-------
2x/(x+2) -4 = 6/x
------
Multiply thru by x(x+2) to get:
2x^2 - 4x(x+2) = 6(x+2)
2x^2 - 4x^2 - 8 = 6x + 12
-2x^2 - 6x -20 = 0
x^2 + 3x + 10 = 0
Use the Quadratic formula to solve for "x":
Cheers,
Stan H.
======================